# Generalized Petermann factor of non-Hermitian systems at exceptional points

### v1

@@ IV.1 Setting the scene
[…]
@@
[…]
@@
×$\times\langle R_{0}|\hat{\Gamma}|R_{0}\rangle\frac{1}{\omega-\Omega_{0}-i(\gamma-\Gamma_{0})/2}.$ (56)
@@
[…]
while nonlinear sufficiently far above the laser threshold feedback reduces this again by a factor of 2, to
[…]
We see that the Petermann factor arises from the matrix elements (57) of the anti-Hermitian part $i\hat{\Gamma}$ of the effective cold-cavity Hamiltonian, which characterizes the openness of the system. On the one hand, this reflects the nature of the physical problem, where the strength of the quantum noise depends on the coupling to external vacuum fluctuations. From this physical perspective, the excess noise can then be attributed to crosstalk with nonlasing modes that overlap with the lasing mode.
@@ IV.2 Generalization to exceptional points
We now evaluate the quantum-limited excess noise near an EP of order $n_{l}$, where we again denote the frequency $\omega_{l}=\Omega_{l}-i\Gamma_{l}/2$, and the right and left eigenstates $|R_{l}\rangle$, $\langle L_{l}|$. The starting point is the general expression Eq. (55), into which we now introduce the approximation (31). Utilizing Eqs. (44), (45), and (57), we obtain the resonant ASE emission spectrum
@@
It is emphasized that neither the intensity in Eq. (62) nor the width in Eq. (63) diverges to infinity at the EP. This recovers the behavior from specific model scenarios [44], generalizes this to EPs or any order, and reexpresses this in terms of the general quantities studied in the present work. The key point is that the excess noise intensity at the EP depends on the spectral strength $\xi_{l}^{2}$. Therefore, the EP Petermann factors developed in the previous section directly carry over to this setting. This extends the general picture that quantum noise is enhanced by mode non-orthogonality to systems operating at EPs. In particular, just as in the original setting of isolated resonances, the EP Petermann factors naturally quantify the noise crosstalk with other modes that overlap with the lasing mode; however, by employing the notions of a right, left, and overall Petermann factor we can now identify and analyze physically and geometric distinct contributions from the right and left eigenmodes.
@@ V Detailed evaluation
with $(N-n)$-dimensional vector $|\psi_{\text{a}}\rangle$ and $n$-dimensional vector $|\psi_{\text{b}}\rangle$. To avoid over-burdening the notation, we focus in this section on a specific EP with eigenvalue $E_{l}\equiv\lambda$ and also drop the index $l$ from the eigenvectors and the spectral strength. In the above two-component representation a right eigenvector of $H$ is
@@
From the dimensionality-reduced eigenvalue equation (71) all eigenvalues of the full Hamiltonian (64) can be retrieved because of the nonlinearity with respect to the eigenvalues. The nonlinear eigenvalue equation (71) is therefore also called isospectral reduction, see, e.g., Ref. [51]. When an eigenvalue $\lambda$ and eigenvector $|r_{\text{b}}\rangle$ of the reduced system is determined the corresponding eigenvector of the full Hamiltonian $H$ can be computed with the help of Eq. (70) as
@@
With a Schur decomposition, we can always bring an Hamiltonian by a unitary transformation into the form in Eq. (64) with $W=0$ or $V=0$. In both cases, the effective Hamiltonian $h(E)$ equals $H_{\text{b}}$. In the following, we consider the situation, where an EPn of the $n\times n$ matrix $H_{\text{b}}$ is also an EPn of the full system.
@@
$N^{n-1}=\xi\frac{|R_{\text{EP}}\rangle\langle L_{\text{EP}}|}{||{R_{\text{EP}}}||_{2}||{L_{\text{EP}}}||_{2}}.$ (81)
@@
$G(E)\sim\frac{\xi}{(E-E_{\text{EP}})^{n}}\frac{|R_{\text{EP}}\rangle\langle L_{\text{EP}}|}{||{R_{\text{EP}}}||_{2}||{L_{\text{EP}}}||_{2}}.$ (82)
@@
$G_{\text{b}}(E)\sim\frac{\xi}{(E-E_{\text{EP}})^{n}}\frac{|r_{\text{EP,b}}\rangle\langle l_{\text{EP,b}}|}{||{R_{\text{EP}}}||_{2}||{L_{\text{EP}}}||_{2}}.$ (83)
@@
$G_{\text{b}}(E)\sim\frac{\xi^{(R,L)}}{(E-E_{\text{EP}})^{n}}\frac{|r_{\text{EP,b}}\rangle\langle l_{\text{EP,b}}|}{||{r_{\text{EP,b}}}||_{2}||{l_{\text{EP,b}}}||_{2}}.$ (84)
@@
$\xi=\xi^{(R,L)}\frac{||{R_{\text{EP}}}||_{2}||{L_{\text{EP}}}||_{2}}{||{p_{\text{b}}R_{\text{EP}}}||_{2}||{p_{\text{b}}L_{\text{EP}}}||_{2}}.$ (85)
@@
$K^{(R,L)}=\left(\frac{\xi}{\xi^{(R,L)}}\right)^{2}=\left(\frac{||{R_{\text{EP}}}||_{2}||{L_{\text{EP}}}||_{2}}{||{p_{\text{b}}R_{\text{EP}}}||_{2}||{p_{\text{b}}L_{\text{EP}}}||_{2}}\right)^{2}.$ (86)
For the Hamiltonian $H$ in Eq. (64) with the eigenvectors in Eqs. (73) and (74) follows
$K^{(R,L)}$ $=$ $\left(1+\langle r_{\text{EP,b}}|V^{\dagger}G_{\text{a}}^{\dagger}(E_{\text{EP}})G_{\text{a}}(E_{\text{EP}})V|r_{\text{EP,b}}\rangle\right)\times$ (87)
$\left(1+\langle l_{\text{EP,b}}|WG_{\text{a}}(E_{\text{EP}})G_{\text{a}}^{\dagger}(E_{\text{EP}})W^{\dagger}|l_{\text{EP,b}}\rangle\right)$
where the vectors $|r_{\text{EP,b}}\rangle$ and $|l_{\text{EP,b}}\rangle$ are normalized to unity. Note that in general $\langle l_{\text{EP,b}}|r_{\text{EP,b}}\rangle\neq 0$ even if $\langle L_{\text{EP}}|R_{\text{EP}}\rangle=0$. This remarkable fact indicates that $|r_{\text{EP,b}}\rangle$ and $|l_{\text{EP,b}}\rangle$ are not the right and left EP-eigenvectors of $H_{\text{b}}$. Note that Eq. (87) is only valid if either $V=0$ or $W=0$. From Eq. (87) or (86) we can confirm that $K^{(R,L)}\geq 1$.
@@
$|E_{j}-E_{\text{EP}}|^{n}\leq\varepsilon||{H_{1}}||_{\text{2}}\,\xi\ .$ (88)
The new quantity $\xi^{(R,L)}$ on the other hand describes the spectral response to perturbations of the subsystem b alone, expressed by the matrix $h_{1}=p_{\text{b}}H_{1}p_{\text{b}}$,
$|E_{j}-E_{\text{EP}}|^{n}\leq\varepsilon||{h_{1}}||_{\text{2}}\,\xi^{(R,L)}\ .$ (89)
In the same spirit, $\xi$ describes the steady-state intensity response of the full system to arbitrary harmonic excitation with strength $P$ and frequency $\omega$ [36]
$||{\psi}||_{2}\leq P\,\frac{1}{|\hbar\omega-E_{\text{EP}}|^{n}}\,\xi\ .$ (90)
@@
$||{p_{\text{b}}\psi}||_{2}\leq P\,\frac{1}{|\hbar\omega-E_{\text{EP}}|^{n}}\,\xi^{(R,L)}\ .$ (91)
@@ V.1 Toy model
To illustrate our results we consider the following $3\times 3$ toy-model Hamiltonian (see Ref. [43])
@@
The response strength of $H_{\text{b}}$ at the second-order EP with eigenvalue $E_{\text{EP}}=E_{\text{b}}$ is $|B|$, see Ref. [36]. We identify this here with the internal response strength, that is, $\xi^{(L)}=|B|$. With $|r_{\text{EP,b}}\rangle=(1,0)^{\text{T}}$, $|l_{\text{EP,b}}\rangle=(0,1)$, and $G_{\text{a}}(E)=(E-E_{\text{a}})^{-1}$ we obtain in a short calculation from Eq. (87) for the EP Petermann factor
@@
As expected, $K^{(R)}$ and $K^{(L)}$ diverge if $E_{\text{a}}=E_{\text{b}}$, which results in a higher-order degeneracy (EP3). This is unless $A=0$, which results in a degeneracy of algebraic multiplicity 3 and geometric multiplicity 2 (instead of 1 for a generic EP), as covered in the recent article [42].
The expression for $K^{(R)}$ is more complicated than $K^{(L)}$ because the eigenvalues in question are placed in the last block of the Schur decomposition. This is as expected, as the decomposition involves a Gram-Schmidt orthogonalisation that starts with the other eigenvectors in the system. Simpler expressions would be obtained by reordering the procedure to give a diagonal $E_{\text{b}},E_{\text{b}},E_{\text{a}}$, or by bringing the system into lower-triangular form. In particular, in our setting, the matrix
@@ VI Realistic example: Two waveguide-coupled microrings
Having discussed a toy model in the previous section we now turn our attention to a more realistic example illustrated in Fig. 1, consisting of two microring cavities coupled to a waveguide; see also Ref. [49]. We use the traveling-wave basis with the counterclockwise propagating wave (CCW) in microring a, the clockwise propagating wave (CW) in microring a, the CCW propagating wave in microring b and the CW propagating wave in microring b. Following coupled-mode theory this system is described by the effective Hamiltonian
@@
with the frequency $\omega_{\text{a}}$ ($\omega_{\text{b}}$ ) of microring a (b), reflectivity $r_{\text{a}}$ of mirror a and reflectivity $r_{\text{b}}$ and transmission coefficient $t$ of mirror b. This system is of the type of Eq. (64), but with $W\neq 0\neq V$. Hence, Eq. (87) does not apply to this more general system.
@@
with the second-order polynom
@@
The software package JCMsuite based on the finite-element method (FEM) [52] is used to solve Eq. (111) for long-lived modes with $\Omega\approx 8.31$. The numerical result are four modes $\psi_{1-4}$ with complex frequencies $\Omega_{1-4}$. However, as shown in Fig. 2(a) two of them have (almost) the same complex frequency $\Omega_{2}=\Omega_{3}$ and mode patterns $|\psi_{2}|=|\psi_{3}|$ indicating an EP of second order. We therefore refer to these two modes as $\Omega_{\text{EP}}$ and $\psi_{\text{EP}}$.
@@
As can be seen in Fig. 3 the EP mode $\psi_{\text{EP}}$ corresponds to a pure CCW wave in cavity b. The two non-EP modes $\psi_{1}$ and $\psi_{4}$ exhibit a field in both cavities; see also Fig. 2(b). Thus, the overlap integral
@@
with $M$ being the area of the two microrings, is finite for all combinations $i,j=1,...,4$. In particular, modes $\psi_{1}$ and $\psi_{4}$ both have an overlap with the EP mode $\psi_{\text{EP}}$ that is $|S|\approx 0.66$ while the overlap between mode $\psi_{1}$ and $\psi_{4}$ is $|S|\approx 0.29$. In addition, the mode $\psi_{\text{EP}}$ shows the self-orthogonal feature with $|S[\psi_{\text{EP}}^{*},\psi_{\text{EP}}]|\approx 6.9\times 10^{-4}$.
@@ VII Conclusion
As a key step to identify the EP Petermann factor, we introduced the concept of internal spectral response strengths that quantify the spectral response to subsystem perturbations and the intensity response to subsystem excitations. An important observation is that the EP subsystem can be constructed in two natural, generally inequivalent, ways, which relate to the generalized left and right eigenspaces associated with the EP. In the setting of dynamical response, the left eigenspace is known to relate to the source of the excitations, while the right eigenspace refers to the detection of the response. Practically, the corresponding EP Petermann factors then capture the difference in the spectral response of the total system relative to the response of the EP subsystem. Therefore, these left and right EP Petermann factors could be determined experimentally by comparing the response in the corresponding source and detection configurations. Finally, these two Petermann factors can be combined, via their geometric mean, into an overall Petermann factor, which itself again carries a well-defined geometric meaning.
@@ Appendix A Derivation of the geometric expressions for the EP Petermann factor
$\hat{N}_{l}^{(R)}\equiv\hat{P}_{l}^{(R)}\hat{N}_{l}\hat{P}_{l}^{(R)}$ $=\sum_{km}A_{k+1,m}|J^{(l)}_{k}\rangle\langle J^{(l)}_{m}|,$ (113)
$\hat{N}_{l}^{(L)}\equiv\hat{P}_{l}^{(L)}\hat{N}_{l}\hat{P}_{l}^{(L)}$ $=\sum_{km}\tilde{A}_{k,m-1}|\tilde{J}^{(l)}_{k}\rangle\langle\tilde{J}^{(l)}_{m}|,$ (114)
@@
$(\hat{N}_{l}^{(R)})^{r}$ $=\sum_{km}A_{k+r,m}|J^{(l)}_{k}\rangle\langle J^{(l)}_{m}|,$ (115)
$(\hat{N}_{l}^{(L)})^{r}$ $=\sum_{km}\tilde{A}_{k,m-r}|\tilde{J}^{(l)}_{k}\rangle\langle\tilde{J}^{(l)}_{m}|.$ (116)

### v2

@@ IV.1 Setting the scene
arising from an isolated cold-cavity mode with complex frequency $\omega_{0}=\Omega_{0}-i\Gamma_{0}/2$ is a Lorentzian of a pump-strength dependent total intensity $I=\int I(\omega)d\omega$ and linewidth $\Delta\omega$, which are related by
$\Delta\omega=\frac{\Gamma_{0}^{2}}{2I}\equiv\Delta\omega_{ST}\quad\mbox{(lasing case)}$ (56)
@@
$\Delta\omega=\frac{\Gamma_{0}^{2}}{I}=2\Delta\omega_{ST}\quad\mbox{(ASE case)},$ (57)
@@
$\times\langle R_{0}|\hat{\Gamma}|R_{0}\rangle\frac{1}{\omega-\Omega_{0}-i(\gamma-\Gamma_{0})/2}.$ (59)
@@
$\Delta\omega=2K\Delta\omega_{ST}\quad\mbox{(ASE case)},$ (63)
while sufficiently far above the laser threshold feedback reduces this again by a factor of 2, to
$\Delta\omega=K\Delta\omega_{ST}\quad\mbox{(lasing case)}.$ (64)
We see that the Petermann factor arises from the matrix elements (60) of the anti-Hermitian part $i\hat{\Gamma}$ of the effective cold-cavity Hamiltonian, which characterizes the openness of the system. On the one hand, this reflects the nature of the physical problem, where the strength of the quantum noise depends on the coupling to external vacuum fluctuations. From this physical perspective, the excess noise can then be attributed to crosstalk with nonlasing modes that overlap with the lasing mode. We note that this quantum limit applies generally also to disordered or wave-chaotic systems [19, 20], while experimental considerations to approach this limit can be found in Refs. [51, 52].
@@ IV.2 Generalization to exceptional points
We now evaluate the quantum-limited excess noise near an EP of order $n_{l}$, where we again denote the frequency $\omega_{l}=\Omega_{l}-i\Gamma_{l}/2$, and the right and left eigenstates $|R_{l}\rangle$, $\langle L_{l}|$. The starting point is the general expression in Eq. (58), into which we now introduce the approximation (31). Utilizing Eqs. (44), (45), and (60), we obtain the resonant ASE emission spectrum
@@
It is emphasized that neither the intensity in Eq. (65) nor the width in Eq. (66) diverges to infinity at the EP. This recovers the behavior from specific model scenarios [44], generalizes this to EPs or any order, and reexpresses this in terms of the general quantities studied in the present work. The key point is that the excess noise intensity at the EP depends on the spectral strength $\xi_{l}^{2}$. Therefore, the EP Petermann factors developed in the previous section directly carry over to this setting. In particular, these EP Petermann factors can in principle be isolated by suitable design of the source (now given by the gain medium) or detector, following the same prescription as in Eqs. (49) and (50). This extends the general picture that quantum noise is enhanced by mode non-orthogonality to systems operating at EPs. In particular, just as in the original setting of isolated resonances, the EP Petermann factors naturally quantify the noise crosstalk with other modes that overlap with the lasing mode; however, by employing the notions of a right, left, and overall Petermann factor we can now identify and analyze physically and geometric distinct contributions from the right and left eigenmodes.
@@ V Detailed evaluation
with (Nn)(N-n)-dimensional vector $|\psi_{\text{a}}\rangle$ and $n$-dimensional vector $|\psi_{\text{b}}\rangle$. To avoid over-burdening the notation, we focus in this section on a specific EP with eigenvalue $E_{l}\equiv\lambda$ and also drop the index $l$ from the eigenvectors and the spectral strength. In the above two-component representation a right eigenvector of $H$ is
@@
From the dimensionality-reduced eigenvalue equation (74) all eigenvalues of the full Hamiltonian (67) can be retrieved because of the nonlinearity with respect to the eigenvalues. The nonlinear eigenvalue equation (74) is therefore also called isospectral reduction, see, e.g., Ref. [53]. When an eigenvalue $\lambda$ and eigenvector $|r_{\text{b}}\rangle$ of the reduced system is determined, the corresponding eigenvector of the full Hamiltonian $H$ can be computed with the help of Eq. (73) as
@@
With a Schur decomposition, we can always bring a Hamiltonian by a unitary transformation into the form in Eq. (67) with $W=0$ or $V=0$. In both cases, the effective Hamiltonian $h(E)$ equals $H_{\text{b}}$. In the following, we consider the situation, where an EPn of the $n\times n$ matrix $H_{\text{b}}$ is also an EPn of the full system.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
For the Hamiltonian $H$ in Eq. (67) with the eigenvectors in Eqs. (76) and (77), it follows for $V=0$
$K^{(R)}=\left(1+\langle l_{\text{EP,b}}|WG_{\text{a}}(E_{\text{EP}})G_{\text{a}}^{\dagger}(E_{\text{EP}})W^{\dagger}|l_{\text{EP,b}}\rangle\right)$ (90)
and for $W=0$
$K^{(L)}=\left(1+\langle r_{\text{EP,b}}|V^{\dagger}G_{\text{a}}^{\dagger}(E_{\text{EP}})G_{\text{a}}(E_{\text{EP}})V|r_{\text{EP,b}}\rangle\right)$ (91)
where the vectors $|r_{\text{EP,b}}\rangle$ and $|l_{\text{EP,b}}\rangle$ are normalized to unity. Note that in general $\langle l_{\text{EP,b}}|r_{\text{EP,b}}\rangle\neq 0$ even if $\langle L_{\text{EP}}|R_{\text{EP}}\rangle=0$. This remarkable fact indicates that $|r_{\text{EP,b}}\rangle$ and $|l_{\text{EP,b}}\rangle$ are not the right and left EP-eigenvectors of $H_{\text{b}}$. From Eqs. (90)-(91) or (89) we can confirm that $K^{(R,L)}\geq 1$.
@@
[…]
The quantity $\xi^{(R,L)}$ on the other hand describes the spectral response to perturbations of the subsystem b alone, expressed by the matrix $h_{1}=p_{\text{b}}H_{1}p_{\text{b}}$,
[…]
In the same spirit, $\xi$ describes the steady-state intensity response of the full system to arbitrary harmonic excitation with strength $P>0$ and frequency $\omega$ [36]
${\psi}||_{2}\leq P\,\frac{1}{|\hbar\omega-E_{\text{EP}}|^{n}}\,\xi.$ (94)
@@
${p_{\text{b}}\psi}||_{2}\leq P\,\frac{1}{|\hbar\omega-E_{\text{EP}}|^{n}}\,\xi^{(R,L)}\ .$ (95)
@@ V.1 Toy model
To illustrate our results, we consider the following $3\times 3$ toy-model Hamiltonian (see Ref. [43])
@@
The response strength of $H_{\text{b}}$ at the second-order EP with eigenvalue $E_{\text{EP}}=E_{\text{b}}$ is $|B|$, see Ref. [36]. We identify this here with the internal response strength, that is, $\xi^{(L)}=|B|$. With $|r_{\text{EP,b}}\rangle=(1,0)^{\text{T}}$, $\langle l_{\text{EP,b}}|=(0,1)$, and $G_{\text{a}}(E)=(E-E_{\text{a}})^{-1}$ we obtain in a short calculation from Eq. (91) for the EP Petermann factor
@@
As expected, $K^{(R)}$ and $K^{(L)}$ diverge if $E_{\text{a}}=E_{\text{b}}$, which results in a higher-order degeneracy, an EP3 (the physical meaning of such a divergence is discussed in Ref. [43]). This is unless $A=0$, which results in a degeneracy of algebraic multiplicity 3 and geometric multiplicity 2 (instead of 1 for a generic EP), as covered in the recent article [42].
The expression for $K^{(R)}$ is more complicated than $K^{(L)}$ because the eigenvalues in question are placed in the last block of the Schur decomposition. This is as expected, as the decomposition involves a Gram-Schmidt orthogonalization that starts with the other eigenvectors in the system. Simpler expressions would be obtained by reordering the procedure to give a diagonal $E_{\text{b}},E_{\text{b}},E_{\text{a}}$, or by bringing the system into lower-triangular form. In particular, in our setting, the matrix
@@ VI Realistic example: Two waveguide-coupled microrings
Having discussed a toy model in the previous section we now turn our attention to a more realistic example illustrated in Fig. 1, consisting of two microring cavities coupled to a waveguide, which contains two partial mirrors. Implementing second-order EPs in such microring-waveguide systems has been first proposed in Ref. [54] and shortly after experimentally realized in Ref. [47]. The extension to higher-order EPs can be found in Ref. [49]. We use the traveling-wave basis with the counterclockwise propagating wave (CCW) in microring a, the clockwise propagating wave (CW) in microring a, the CCW propagating wave in microring b and the CW propagating wave in microring b. Following coupled-mode theory this system is described by the effective Hamiltonian
@@
with the frequency $\omega_{\text{a}}$ ($\omega_{\text{b}}$ ) of microring a (b), reflectivity $r_{\text{a}}$ of mirror a and reflectivity $r_{\text{b}}$ and transmission coefficient $t$ of mirror b. This system is of the type of Eq. (67), but with $W\neq 0\neq V$. Hence, Eqs. (90) and (91) do not apply to this more general system.
@@
with the second-order polynomial
@@
The software package JCMsuite based on the finite-element method (FEM) [55] is used to solve Eq. (115) for long-lived modes with $\Omega\approx 8.31$. The numerical results are four modes $\psi_{1-4}$ with complex frequencies $\Omega_{1-4}$. However, as shown in Fig. 2(a) two of them have (almost) the same complex frequency $\Omega_{2}=\Omega_{3}$ and mode patterns $|\psi_{2}|=|\psi_{3}|$ indicating an EP of second order. We therefore refer to these two modes as $\Omega_{\text{EP}}$ and $\psi_{\text{EP}}$.
@@
As can be seen in Fig. 3, the EP mode $\psi_{\text{EP}}$ corresponds to a pure CCW wave in cavity b. The two non-EP modes $\psi_{1}$ and $\psi_{4}$ exhibit a field in both cavities; see also Fig. 2(b). Thus, the overlap integral
@@
with $M$ being the area of the two microrings, is finite for all combinations i,j=1,,4i,j=1,...,4. In particular, modes $\psi_{1}$ and $\psi_{4}$ both have an overlap with the EP mode $\psi_{\text{EP}}$ that is $|S|\approx 0.66$ while the overlap between mode $\psi_{1}$ and $\psi_{4}$ is $|S|\approx 0.29$. In addition, the mode $\psi_{\text{EP}}$ shows the self-orthogonal feature with $|S[\psi_{\text{EP}}^{*},\psi_{\text{EP}}]|\approx 6.9\times 10^{-4}$.
@@ VII Conclusion
As a key step to identify the EP Petermann factor, we introduced the concept of internal spectral response strengths that quantify the spectral response to subsystem perturbations and the intensity response to subsystem excitations. An important observation is that the EP subsystem can be constructed in two natural, generally inequivalent, ways, which relate to the generalized left and right eigenspaces associated with the EP. In the setting of dynamical response, the left eigenspace is known to relate to the source of the excitations, while the right eigenspace refers to the detection of the response. Practically, the corresponding EP Petermann factors then capture the difference in the spectral response of the total system relative to the response of the EP subsystem. Therefore, these left and right EP Petermann factors could be determined experimentally by comparing the total response power (48) with the response powers (49) and (50) in the corresponding source and detection configurations. First experimental steps into determining the left and right eigenstates have been reported in Ref. [57]. Finally, these two Petermann factors can be combined, via their geometric mean, into an overall Petermann factor, which itself again carries a well-defined geometric meaning.
@@ Appendix A Derivation of the geometric expressions for the EP Petermann factor
[…]
@@
[…]